# Faithfulness-Aware Uncertainty Quantification for Fact-Checking the Output of Retrieval Augmented Generation

### v2

@@ 4.3 Short-Form QA Results
For short-form QA, we compute three metrics (ROC-AUC, PR-AUC, and PRR) for each UQ method across two models (Llama 3B and Falcon 3B) and four QA datasets. For each metric and dataset, we also determine the relative performance rank of each UQ method (1st best, 2nd best, etc.). Since these detailed results vary across datasets, we report only the mean metric values and mean ranks (averaged over the four datasets) to provide a clearer and more concise summary, following the approach of Vashurin et al. (2025). These summary results are shown in Table 3. Full per-dataset results are provided in Appendix A.
For Llama 3B, the condition-calibrated franq version outperforms all baselines, achieving the best mean rank and mean value across all three metrics. The calibrated franq variant ranks second, showing the top mean rank for ROC-AUC and second-best for PRR, as well as second-best mean value for both ROC-AUC and PRR. Among unsupervised methods, DegMat shows strong performance, ranking second in PR-AUC under both mean rank and mean value.
For Falcon 3B, the condition-calibrated franq method again leads overall, achieving the top mean rank across all metrics, and the best PR-AUC and PRR in mean value, while ranking second in ROC-AUC. XGBoost, trained on all UQ features, ranks second overall, placing second in ROC-AUC and PR-AUC under mean rank, and first in ROC-AUC under mean value. DegMat remains competitive for Falcon 3B as well, with the best PR-AUC and second-best ROC-AUC and PRR in mean rank, and second-best PRR in mean value.
@@ 4.4 Ablation Studies
We conducted ablation studies to evaluate the contributions of individual franq components, the effect of training set size on supervised variants, and the calibration quality of the tested UQ methods. This section summarizes the key findings, and a more detailed analysis is provided in Appendix D.
Analysis of franq components. We begin by analyzing different combinations of $\text{UQ}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$ within condition-calibrated franq (see Figure 11 in in Appendix D). For long-form QA with Llama 3B, performance is highly sensitive to the choice of $\text{UQ}_{\text{unfaith}}$, with Parametric Knowledge emerging as the most effective option. In contrast, $\text{UQ}_{\text{unfaith}}$ has a relatively minor impact. For short-form QA with Llama 3B, franq is robust across multiple UQ configurations, suggesting greater flexibility in component selection.
Next, we evaluate alternative faithfulness estimation strategies (see Tables 12 and 13). Substituting AlignScore probabilities with binary thresholding consistently degrades performance, underscoring the benefit of using continuous faithfulness signals. Surprisingly, calibrating AlignScore with gold-standard labels also reduces performance, indicating that the raw AlignScore values are more effective when used directly within franq.
Impact of the training set size. We also examine the effect of the training set size on both franq and XGBoost variants (see Figure 12). The unsupervised franq baseline remains stable, while supervised versions with calibration generally improve with more data. In long-form QA, condition-calibrated franq peaks at 300 training instances before slightly declining; for short-form QA, performance saturates around 120 instances. Across all training sizes, supervised franq methods consistently outperform XGBoost ones.
XGBoost approximates the decision-making behavior of franq. To better understand XGBoost’s behavior, we analyze its first decision tree trained on franq features (see Figure 13 in Appendix D). We find that the tree mimics franq’s decision logic: initially splitting on AlignScore, followed by MaxNLI or Parametric Knowledge depending on the score. This suggests that supervised methods like XGBoost learn decision patterns closely aligned with those encoded in franq.
Calibration evaluation. Finally, we assess the calibration of all UQ methods using Expected Calibration Error (ECE; Guo et al. (2017)); see Table 14 in Appendix D. As expected, the two calibrated franq variants achieve the lowest ECE scores for both long- and short-form QA, indicating that they produce well-calibrated confidence estimates aligned with empirical accuracy.
@@ 5 Related Work
5.1 Uncertainty Quantification for RAG
Several uncertainty quantification methods have been proposed for RAG that take into account the influence of retrieved knowledge on the LLM output. Chuang et al. (2024) detect hallucinations with a linear classifier that leverages the lookback ratio, the ratio of attention weights on the context versus newly generated tokens. The method of Sun et al. (2025) uses regression on features measuring the model’s reliance on external extracted knowledge (via attention scores) and internal parametric knowledge (from its own weights). Another UQ method for RAG gets an uncertainty score that estimates the similarity between text fragments through the signal-to-noise ratio of the output probabilities of the model sample (Li et al., 2024). The hallucination detection method for RAG proposed by Hu et al. (2024) estimates the correlation between the model’s input and output by computing a relevance matrix between the prompt (including retrieved context) and the response, and then it is classified as hallucinations.
All of the aforementioned methods identify hallucinations with respect to the retrieved context, but do not assess them in relation to world knowledge, which represents a notable shortcoming. Moreover, these approaches incur additional computational overhead due to the need for an auxiliary model and its associated training process.
When RAG does not incorporate retrieved context during generation, various methods estimate uncertainty solely based on the model’s internal knowledge, without accounting for retrieval-related uncertainty. These include both white-box approaches (Fomicheva et al., 2020; Kadavath et al., 2022; Kuhn et al., 2023; Fadeeva et al., 2024; Duan et al., 2024) and black-box approaches (Fomicheva et al., 2020; Lin et al., 2024).
5.2 Hallucination and Factuality Datasets for RAG
The task of hallucination detection in RAG requires labeled data that captures model hallucinations in the form of corresponding class annotations. RAGTruth (Niu et al., 2024) is a multi-domain dataset for RAG hallucination detection, featuring manually annotated and span-level classified hallucinations. However, the hallucination classes in this dataset do not account for cases where the model generates knowledge independently of the retrieved context and is then verified separately for hallucinations, as such responses may still be factually correct and useful to the user.
Knowledge-based dialogue datasets bear some similarity to RAG datasets, as they also include an external knowledge base to support the conversation. Wizard of Wikipedia Dinan et al. (2019) and FaithDial (Dziri et al., 2022), for example, are structured as user–model dialogues grounded in Wikipedia content. Unlike RAG datasets, however, these datasets place greater emphasis on maintaining coherent dialogue and conversational consistency. Moreover, as in the RAGTruth dataset, any information not directly grounded in the provided context is automatically treated as hallucination, regardless of its factual accuracy.
Hallucination benchmarks for RAG are often constructed using QA datasets, where relevant context is provided alongside the question (Friel et al., 2024; Moskvoretskii et al., 2025). In these datasets, hallucinations are defined strictly with respect to the provided context.
@@ 6 Conclusion and Future Work
In this paper, we introduce franq (Faithfulness-based Retrieval Augmented UNcertainty Quantification), a new method that quantifies the factuality of claims in RAG output based on their faithfulness. Evaluated on both long-form and short-form QA tasks across multiple LLMs, franq outperforms existing unsupervised UQ baselines, RAG-specific methods, and supervised classifiers. We also introduce a new long-form QA dataset annotated for both factuality and faithfulness using a hybrid of automatic and manual labeling.
Our approach opens several promising directions for future research. One is extending uncertainty modeling to the retrieval stage itself, allowing systems to account for noisy, incomplete, or conflicting evidence. Future work could also explore how franq’s uncertainty signals can guide generation-time control or post-editing, enabling more reliable and interpretable RAG systems.
@@
While franq provides strong hallucination detection performance on average, it does not guarantee ideal hallucination detection in every situation, as it is a challenging task.
The method assumes that retrieved evidence is always factual and takes precedence over the LLM’s parametric knowledge. In theory, this can be achieved through careful selection and curation of document sources within the search index. However, ensuring complete factual accuracy in real-world applications might be challenging as the size of the index grows.
Since franq leverages the calibration of its components, it might be considered as supervised. To address this concern, we show that this method also outperforms supervised methods.
@@
franq is designed to reduce the spread of factual errors by enhancing the interpretability and reliability of language model outputs. By distinguishing between factuality and faithfulness, it helps prevent misclassification of factually correct but unsupported claims. However, the system does not actively prevent the generation of hallucinations and instead relies on downstream filtering. Its effectiveness, therefore, depends on integration into larger pipelines with proper safeguards.
franq assumes the retrieved context is factual and trustworthy. In real-world applications, retrieved documents may be biased, outdated, or incorrect, which could compromise the method’s output. Careful curation of retrieval sources and monitoring of retrieval quality are crucial to avoid reinforcing harmful biases or misinformation.
The dataset used in evaluation relies on outputs from GPT-4o. While we manually validate a subset of annotations, some inherent biases from the underlying model may persist. We encourage future work to explore more diverse annotation strategies, including community-sourced validation.
Improving factuality estimation can support safer AI deployment, especially in knowledge-intensive domains such as education, healthcare, or law. However, the system should not be considered a replacement for human fact-checkers. It is best used as a decision-support tool rather than a source of truth.
@@ Appendix A Additional short-form QA results
In Table 3 of the main text, we report aggregated results for short-form QA using MeanRank and MeanValue for ease of presentation. Here, we provide the full results for each of the four QA datasets (Natural Questions, PopQA, TriviaQA, SimpleQA) for both Llama 3B (see Table 4) and Falcon 3B (see Table 5).
For Llama 3B, franq calibrated and franq condition-calibrated methods consistently rank among the top performers. They are the top two methods on TriviaQA and SimpleQA. On PopQA, franq condition-calibrated ranks among the top three methods, alongside Semantic Entropy and Max Token Entropy. On Natural Questions, it ranks in the top four, along with DegreeMatrix, Eccentricity, and Sum of Eigenvalues. Overall, both franq variants achieve the best average performance across all datasets.
@@
Max Sequence Probability .680 .440 .292 .745 .550 .421 .774 .529 .478 .833 .712 .625
@@
Table 4: Results on 4 QA datasets for Llama-3b.
For Falcon 3B, franq condition-calibrated achieves the top performance on TriviaQA and second-best performance on Natural Questions. It also ranks among the top three methods on PopQA and among the top four on SimpleQA, alongside Degree Matrix, Sum of Eigenvalues, and XGBoost (all features). On average, franq condition-calibrated is the leading method across the four datasets.
@@
Max Sequence Probability .599 .555 .186 .653 .649 .259 .590 .487 .163 .625 .820 .416
@@
Table 5: Results on 4 QA datasets for Falcon-3b.
@@ Appendix B Prompts and setup
Contents (not necessarily includes answer to the following question): Title: {title1} Content: {retrieval1} Title: {title2} Content: {retrieval2} Title: {title5} Content: {retrieval5} Question: {question} Answer (single line):
@@
B.1 Short-form QA
For short-form QA experiments, we used each input question along with the top-5 retrieved passages from Wikipedia. The prompt format used across all evaluated models is shown in Figure 3. For annotation, we provided GPT-4 with the input question, the model-generated answer, and the corresponding gold answer, and asked it to categorize the model’s response as correct, incorrect, or not attempted (the latter category was excluded from evaluation). This annotation method follows the approach proposed by Wei et al. (2024). Table 6 presents the statistics of all short-form QA datasets.
@@ B.1 Short-form QA
B.2 Long-form QA
Figure 4: Prompt template used wit GPT-4o for decomposing an answer into a set of atomic claims.
For long-form QA experiments, we used each input question along with the top-3 retrieved passages from Wikipedia. The prompt format employed across all evaluated models is illustrated in Figure 3. We then decomposed the extracted answers into atomic claims using the prompt shown in Figure 4. For each claim, we identified its corresponding span in the original sentence using the prompt in Figure 5. Claims for which we could not extract spans (due to inconsistencies in annotation, e.g., words appearing in a different order than in the text) were excluded from evaluation. Finally, we annotated the resulting claims for both factuality and faithfulness using a combination of automatic annotation and subsequent manual validation (see Appendix B.3 and  B.4). Table 7 presents the statistics of all long-form QA dataset.
@@ B.2 Long-form QA
Llama3b 500 1282 90.5 % 6.2 % 3.3 % 37.3 % 62.6 % 0.1 % 1725.4
Falcon3b 500 1048 91.4 % 6.0 % 2.6 % 38.2 % 61.5 % 0.3 % 1720.2

### v3

@@ 4.3 Short-Form QA Results
For short-form QA, we evaluate UQ methods using PR-AUC and PRR across four models (Llama 3B Instruct, Llama 8B Instruct, Falcon 3B Base, Gemma 12B Instruct) and four datasets. To account for dataset variability, we report mean scores averaged over datasets, following Vashurin et al. (2025) (Table 3); per-dataset results appear in Appendix A.
Condition-calibrated franq achieves the best mean performance across all models and both measures, except for mean PRR on Gemma 12B Instruct, where it ranks second. Calibrated franq ranks second for PRR on Llama 3B Instruct and for both PR-AUC and PRR on Llama 8B Instruct. Among unsupervised methods, Degree Matrix and Lexical Similarity perform strongly, ranking second-best in several settings.
@@ 4.4 Ablation Studies
In this section, we summarize the main observations from ablation studies examining (1) the contribution of franq’s components, (2) robustness to retrieval noise, (3) the effect of supervision and (4) computational efficiency. Complete experimental descriptions, tables, and additional ablations are provided in Appendix D.
(a) PRR on long-form QA dataset.
(b) PRR on short-form QA benchmark (mean across 4 datasets).
Analysis of franq’s components. Figure 2 reports the PRR of franq with condition calibration for different choices of $\text{UQ}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$, evaluated on a subset of 200 questions from each dataset using the Llama 3B Instruct model.
On the long-form QA dataset (see Figure 2(a)), performance is largely insensitive to the choice of $\text{UQ}_{\text{faith}}$, whereas the choice of $\text{UQ}_{\text{unfaith}}$ is decisive: using Parametric Knowledge as $\text{UQ}_{\text{unfaith}}$ yields the best PRR across essentially all $\text{UQ}_{\text{faith}}$ options.
On short-form QA (see Figure 2(b)), many combinations of $\text{UQ}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$ perform similarly, suggesting that franq is relatively robust to these design choices. The configuration used in our short-form experiments (Semantic Entropy for $\text{UQ}_{\text{faith}}$ and Sum of Eigenvalues for $\text{UQ}_{\text{unfaith}}$) achieves best observed value of PRR =0.553=0.553.
We further examine alternative faithfulness modeling strategies in Appendix D.1. Replacing continuous AlignScore probabilities with binary thresholding degrades performance, underscoring the value of probabilistic faithfulness weighting.
Robustness to retrieval noise. We evaluate franq under corrupted retrievals by randomly replacing a fraction of retrieved documents with unrelated passages, simulating noisy RAG settings (Appendix A.1). Results show that even with 50% retrieval noise, calibrated and condition-calibrated franq achieve top-ranked performance, indicating that the method can effectively handle substantial amounts of irrelevant evidence.
Effect of supervision. We analyze the impact of training set size on supervised franq variants by varying the number of labeled calibration examples used for training (see Appendix D.2). Performance improves with additional data and then saturates, with the optimal training size occurring at approximately 300 instances for long-form QA, while for short-form QA performance stabilizes at around only 120 training samples.
Computational efficiency. We analyze the runtime overhead, training cost, and model size of franq’s uncertainty components in Appendix E. Overall, franq introduces minimal training cost (fitting isotonic regression takes less than a second), while producing extremely compact models (on the order of hundreds of bytes), making calibrated franq variants practical for real-world deployment.
@@ 5 Related Work
Uncertainty Quantification for RAG. Several UQ methods for RAG analyze how retrieved knowledge influences LLM outputs, including lookback-ratio classifiers (Chuang et al., 2024), feature-based regression of retrieved versus parametric reliance (Sun et al., 2025), uncertainty estimation via the signal-to-noise ratio of output probabilities across samples (Li et al., 2024), and prompt–response relevance modeling (Hu et al., 2024). These methods evaluate hallucinations only relative to retrieved context and often incur computational overhead, while search-based approaches such as SAFE Wei et al. (2024b) rely on LLM agents and web verification. In contrast, franq is a lightweight, self-contained UQ framework that probabilistically combines faithfulness to retrieved context with truthfulness under both faithful and unfaithful conditions, without additional training or external verification.
When retrieval is absent, uncertainty is typically estimated from internal model knowledge using white-box (Fomicheva et al., 2020; Kadavath et al., 2022; Kuhn et al., 2023; Fadeeva et al., 2024; Duan et al., 2024) or black-box (Fomicheva et al., 2020; Lin et al., 2024) methods. franq unifies these settings by jointly modeling retrieval-related and intrinsic uncertainty within a single probabilistic framework.
Factuality/Hallucination Datasets for RAG. Hallucination detection in RAG relies on labeled datasets of factual errors. RAGTruth (Niu et al., 2024) offers multi-domain, span-level annotations, but excludes cases where models produce correct information independent of the retrieved context.
Knowledge-grounded dialogue datasets such as Wizard of Wikipedia Dinan et al. (2019) and FaithDial (Dziri et al., 2022) pair responses with external sources but prioritize conversational coherence, treating any content not grounded in the provided context as hallucination regardless of factual correctness.
QA-based benchmarks such as RAGBench (Friel et al., 2024) and AdaptiveRAG (Moskvoretskii et al., 2025) define hallucinations strictly relative to the given context, whereas franq estimates factuality even when generation diverges from retrieved evidence, enabling a more comprehensive assessment of model reliability.
@@ 6 Conclusion and Future Work
We introduced franq, a new method for quantifying the factuality of claims in RAG output based on their faithfulness. Across both long-form and short-form QA tasks and multiple LLMs, franq consistently outperforms existing unsupervised UQ baselines, RAG-specific methods, and supervised classifiers. We also presented a new long-form QA dataset annotated for both factuality and faithfulness using a hybrid of automatic and manual labeling.
Our approach opens several promising directions for future research. One direction is to extend uncertainty modeling to the retrieval stage, thus allowing systems to account for noisy, incomplete, or conflicting evidence. Another is to leverage franq’s uncertainty signals for generation-time control and post-editing, thus enabling more reliable and interpretable RAG systems.
@@
While franq provides strong hallucination detection performance on average, it does not guarantee ideal hallucination detection in every situation, as this is a challenging task.
franq assumes that the retrieved evidence is always factual and takes precedence over the LLM’s parametric knowledge. In theory, this can be achieved through careful selection and curation of document sources within the search index. However, ensuring complete factual accuracy in real-world applications might be challenging as the size of the index grows.
Since franq leverages the calibration of its components, it might be considered as supervised. To address this concern, we showed that it also outperforms supervised methods.
@@
franq is designed to reduce the spread of factual errors by enhancing the interpretability and reliability of language model outputs. By distinguishing between factuality and faithfulness, it helps prevent misclassification of factually correct but unsupported claims. However, franq does not actively prevent the generation of hallucinations and instead relies on downstream filtering. Its effectiveness, therefore, depends on integration into larger pipelines with proper safeguards.
franq assumes that the retrieved context is factual and trustworthy. In real-world applications, the retrieved documents may be biased, outdated, or incorrect, which could compromise the method’s output. Careful curation of retrieval sources and monitoring of retrieval quality are crucial to avoid reinforcing harmful biases or misinformation.
The dataset used for evaluation relies on the output from GPT-4o. While we manually validated a subset of the annotations, some inherent biases from the underlying model may persist. We encourage future work to explore more diverse annotation strategies, including community-sourced validation.
Improving the factuality estimation can support safer AI deployment, especially in knowledge-intensive domains such as education, healthcare, or law. However, the system should not be considered a replacement for human fact-checkers. It is best used as a decision-support tool rather than a source of truth.
@@ Appendix A Additional Short-form QA Results
In Table 3 of the main text, we report aggregated results for short-form QA using mean values for ease of presentation. Here, we provide the full results for each of the four QA datasets (Natural Questions, PopQA, TriviaQA, SimpleQA) for both Llama 3B Instruct (see Table 6) and Falcon 3B Base (see Table 7).
For Llama 3B Instruct, franq calibrated and franq condition-calibrated methods consistently rank among the top performers. They are the top two methods on TriviaQA and SimpleQA. On PopQA, franq condition-calibrated ranks among the top three methods, alongside Semantic Entropy and Max Token Entropy. On Natural Questions, it ranks in the top four, along with DegreeMatrix, Eccentricity, and Sum of Eigenvalues. Overall, both franq variants achieve the best average performance across all datasets.
For Falcon 3B Base, franq condition-calibrated achieves the top performance on TriviaQA and second-best performance on Natural Questions. It also ranks among the top three methods on PopQA and among the top four on SimpleQA, alongside Degree Matrix, Sum of Eigenvalues, and XGBoost (all features). On average, franq condition-calibrated is the leading method across the four datasets.
Method MeanValue $\uparrow$
PRAUC $\uparrow$ PRR $\uparrow$
General Baselines
Max Sequence Prob. .638 .489
Mean Token Entropy .651 .460
CCP .629 .472
Lexical Similarity .668 .512
Degree Matrix .687 .548
Sum of Eigenvalues .686 .536
Semantic Entropy .666 .537
SentenceSAR .648 .523
RAG-specific Baselines
AlignScore .507 .234
Parametric Knowledge .502 .220
XGBoost
XGBoost (all UQ features) .684 .544
XGBoost (FRANQ features) .579 .402
franq
FRANQ no calibration .586 .393
FRANQ calibrated .692 .549
FRANQ condition-calibrated .695 .553
Table 4: Aggregated results with shuffled retrievals on 4 QA datasets for Llama 3B Instruct.
Method NQ PopQA TriviaQA SimpleQA Mean Value
PRAUC $\uparrow$ PRR $\uparrow$ PRAUC $\uparrow$ PRR $\uparrow$ PRAUC $\uparrow$ PRR $\uparrow$ PRAUC $\uparrow$ PRR $\uparrow$ PRAUC $\uparrow$ PRR $\uparrow$
General Baselines
Max Sequence Prob. .477 .216 .645 .509 .556 .477 .874 .755 .638 .489
Mean Token Entropy .571 .346 .678 .510 .580 .460 .777 .525 .652 .460
CCP .494 .269 .629 .481 .571 .482 .822 .657 .629 .472
Lexical Similarity .580 .370 .696 .547 .579 .489 .819 .639 .669 .511
Degree Matrix .570 .359 .661 .537 .645 .553 .871 .745 .687 .549
Sum of Eigenvalues .573 .360 .664 .516 .637 .542 .872 .728 .687 .537
Semantic Entropy .537 .317 .697 .587 .565 .493 .867 .752 .667 .537
SentenceSAR .465 .217 .695 .546 .566 .517 .866 .811 .648 .523
RAG-specific Baselines
AlignScore .493 .259 .506 .225 .417 .213 .613 .240 .507 .234
Parametric Knowledge .406 .104 .564 .342 .494 .377 .545 .059 .502 .221
XGBoost
XGBoost (all UQ features) .561 .354 .657 .474 .656 .594 .860 .753 .684 .544
XGBoost (FRANQ features) .488 .269 .579 .376 .471 .370 .779 .592 .579 .402
franq
FRANQ no calibration .419 .127 .626 .452 .527 .417 .774 .575 .587 .393
FRANQ calibrated .601 .401 .675 .549 .624 .529 .868 .717 .692 .549
FRANQ condition-calibrated .603 .413 .675 .549 .636 .536 .867 .715 .695 .553
[…]
@@
General Baselines
Max Sequence Prob. .680 .440 .292 .745 .550 .421 .774 .529 .478 .833 .712 .625
Mean Token Entropy .723 .503 .389 .768 .607 .455 .796 .569 .523 .809 .697 .555
@@
Table 6: Results on 4 QA datasets for Llama 3B Instruct.
@@
General Baselines
Max Sequence Prob. .599 .555 .186 .653 .649 .259 .590 .487 .163 .625 .820 .416
Mean Token Entropy .599 .542 .184 .657 .662 .279 .557 .432 .108 .656 .814 .396
@@
Table 7: Results on 4 QA datasets for Falcon 3B Base.
A.1 Short-form QA with Corrupted Retrievals
Real-world retrieval systems are often imperfect, occasionally introducing irrelevant passages that degrade both LLM and FRANQ performance. To simulate this, we randomly shuffled 50% of the retrieved passages across four QA datasets, ensuring no shuffled sample retained its original retrievals. Corrupted retrievals added substantial noise to the prompt, causing an average 7% accuracy drop. Nevertheless, FRANQ methods remained robust, maintaining competitive results (aggregated in Table 4, full results in Table 5). Under this setting, FRANQ condition-calibrated and FRANQ calibrated variants achieve top-1 and top-2 overall performance in average quality across the four datasets, measured by PR-AUC and PRR.
@@ Appendix B Prompts and Setup
B.1 Short-form QA
For short-form QA experiments, we paired each question with the top-5 retrieved Wikipedia passages and used the prompt format in Figure 4. For annotation, GPT-4o was given the question, model-generated answer, and gold answer, and asked to label responses as correct, incorrect, or not attempted (excluded from evaluation), following Wei et al. (2024a). Table 8 reports dataset statistics.
B.2 Long-form QA
For long-form QA experiments, we used each question with the top-3 retrieved Wikipedia passages. All models followed the prompt format shown in Figure 4. Extracted answers were decomposed into atomic claims using the prompt in Figure 6, and each claim was matched to its corresponding span in the original sentence using Figure 6. Claims without identifiable spans (e.g., due to annotation inconsistencies) were excluded. The remaining claims were annotated for factuality and faithfulness using automatic annotation followed by manual validation (Appendix B.3, B.4). Table 9 reports dataset statistics.
Compared to prior decomposition methods such as FActScore Min et al. (2023), our approach is more careful: we decompose entire texts rather than individual sentences to reduce redundancy and ambiguity, and we produce decontextualized claims to simplify verification. Claim quality was further examined during manual validation in complex cases.
Contents (not necessarily includes answer to the following question): Title: {title1} Content: {retrieval1} Title: {title5} Content: {retrieval5} Question: {question} Answer (single line):
@@ B.2 Long-form QA
Llama 3B Instruct 600 1182 91.0 % 5.8 % 3.1 % 37.3 % 62.6 % 0.1 % 1725.4
Falcon 3B Base 600 948 91.4 % 6.0 % 2.6 % 38.2 % 61.5 % 0.3 % 1720.2
Llama 8B Instruct 300 500 89.4 % 5.0 % 5.6 % 34.6 % 64.6 % 0.8 % 1856.4
Gemma 4B Instruct 300 500 88.8 % 5.7 % 5.5 % 44.7 % 54.5 % 0.8 % 1708.3
@@
Figure 5: Prompt template used with GPT-4o for decomposing an answer into a set of atomic claims.